Circular No. 6552

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997X IN NGC 4691 S. Nakano, Sumoto, Japan, reports the discovery of a supernova (mag

13.6) by Masakatsu Aoki (Tsukioka-cho, Toyama) on nine unfiltered CCD frames taken around Feb. 1.76 UT with a 0.43-m f/6 reflector. Y. Kushida (Yatsugatake South Base Observatory) provides his measurement of SN 1997X (mag 13.5) from an unfiltered CCD image taken by R. Kushida with a 0.40-m f/5 reflector on Feb. 1.816: α = 12^{h}48^{m}14^{s}.28, δ = −3^{o}19′58′′.5 (equinox 2000.0), which is ∼ 7′′.2 east and 0′′.3 north of the center of NGC 4691 (this offset is necessarily uncertain due to the irregular nature of the galaxy’s nucleus). No star was present at the location of SN 1997X on Aoki’s most recent previous patrol films of this galaxy taken on Jan. 16, nor was there anything on Kushida’s most recent patrol frame from Jan. 6.

N. B. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports: “A reduction by A. Clocchiatti, M. M. Phillips, and Suntzeff of a spectrum (resolution 600; range 410–750 nm) obtained by D.-W. Kim, Harvard-Smithsonian Center for Astrophysics (CfA) on Feb. 4.3 UT with the CTIO 1.5-m telescope shows that SN 1997X is probably a type-Ic supernova shortly after maximum light. A strong absorption feature (FWHM 19.4 nm) centered at 566 nm resembles the blend of Na I D and He I 587.6nm in SN 1994I, 10 days after maximum. A weaker absorption centered at ∼ 645 nm might correspond to He I 667.8-nm, indicating that He I lines in this object are stronger than in SN 1994I. Narrow Na I D interstellar lines are weak or absent. R. Covarrubias (CTIO) obtained BVI photometry of the supernova on Feb. 4.3 with the CTIO 0.9-m telescope, yielding preliminary photometry (±0.05): B = 14.75, V = 13.71, I = 12.85. SN 1997X is located 8′′.6 east and 1′′.3 north of the galaxy’s nucleus; note that there is also a bright nuclear knot situated 3′′.2 west and 1′′.2 north of the galaxy’s center, which could be confused with the supernova as it fades.”

P. Garnavich and R. Kirshner, CfA, report that spectra of SN 1997X were obtained on Feb. 4.47 UT with the 1.5-m Tillinghast telescope by P. Berlind. The spectra show a fairly red continuum with deep, broad features indicative of a supernova past maximum. The lack of a prominent Hα feature or deep Si II (615 nm) implies that this is a type-Ib/c supernova. Features with observed minima at 567.5 and 623.0 nm are likely to be Na I and Si II, respectively. The overall spectrum is similar to that of the type-Ic SN 1987M (Filippenko 1992, Ap.J. 384, L37) except that the absorption lines indicate a larger ejection velocity in SN 1997X.

1997 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green